OpenStudy (anonymous):

calculate arcsin 1, without using calculator

OpenStudy (whpalmer4):

\[\arcsin(1) = \theta\]where \[\sin\theta = 1\] What is the first value of \(theta\) you encounter on your unit circle where \(\sin \theta = 1\)?

OpenStudy (anonymous):

90?

OpenStudy (anonymous):

Correct.

OpenStudy (whpalmer4):

as long as that is 90 degrees. just plain 90 is incorrect, as without an indication to the contrary, arguments to the trig functions are assumed to be in radians.

OpenStudy (anonymous):

thank you for your help. i appreciate it